Given that the odds of an event A occurring is 5, what is the risk of event A happening?
1
Expert's answer
2020-10-02T10:48:51-0400


Odds for a means odds in favor of A are 5 and against are 1.


So total outcomes =5+1=6

Hence risk of A = percentage of A =(5/6)*100%= 83.33%